A weighted Gauss quadrature formula for the weight function x is در . ) + - ) ربه ) = ar( 1 )

(a) The formula is exact for f(1) = 1, 1, 2, ..., , what is m ?

(b) Use this formula to approximate the integral (keep 9 decimal places)

1 Jo x sin r de.

(c) Compute the error. (Note: The error is less than 10^{-4} .)
Answer #1

(stroduct;_;(三) ( 34 ) 一) [Q) - 4CM) =|| (一 )x1 + + x = -- 19 十一 - 86 361 1-2 4 → f[n) = X n2 du = 13 3 To | 10 13 ) 『},(1+- 2156 + 3x6 26x50 25x36, + 36 + 21. AX25 200 4425 26x50 + 3x6 25 X36 150 300 30 = 2 2 1 6 + 2 x 2 30 = 2 + 25 = 63 + 12 == 0.3011 606 79. Error 1168 679 - 0.301130765 = 0.000037914 = 0.379 14 x 10-4